OpenStudy (anonymous):

Medal & fan-What is the solution of n2 -49=0? a. -7 b. 7 c.#7 d.No solution my answer b

OpenStudy (anonymous):

Is it \[n ^{2}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then yes, It is B.

OpenStudy (anonymous):

OK TY I have another one -

OpenStudy (anonymous):

Ok ^.^

OpenStudy (shiraz14):

what's C? is it +/-7?

OpenStudy (anonymous):

What is the solution of x2+64=0? a.-5 b.8 c.#8 d.no solution

OpenStudy (anonymous):

It's \[x^{2}\] Right?

OpenStudy (anonymous):

yeap

OpenStudy (shiraz14):

@JuniorEinstein1: What's the # sign for option C in both cases?

OpenStudy (anonymous):

\[\pm both c symbols\]

OpenStudy (shiraz14):

@JuniorEinstein1: OK, thanks. Then the answer to your first question is C.

OpenStudy (anonymous):

It's not A, Or B. So It's either C or D.

OpenStudy (anonymous):

I think it's D though.

OpenStudy (anonymous):

@ shiraz are you sure about example #1

OpenStudy (shiraz14):

@JuniorEinstein1: The answer to your 2nd question is D.

OpenStudy (anonymous):

ty I picked d too But the first one is def c ?

OpenStudy (shiraz14):

For Example #1, n^2 - 49 = (n-7)(n+7) = 0 i.e. n = 7 or n = -7 So the answer is C.

OpenStudy (anonymous):

When you have an exponent you multiply it by itself...@shiraz14

OpenStudy (anonymous):

What is the lenght of a square with an area of 144x2? a.12 b.12x c.#12x d.no solution c ?

OpenStudy (shiraz14):

@blueweek77 : this is a quadratic equation which has 2 roots on the real number line - one at -7 and one at +7. This is not the same as saying that n = square root(49), in which case, n = 7 (since the square root is defined as >0).

OpenStudy (anonymous):

\[\pm for answer c\]

OpenStudy (shiraz14):

@ JuniorEinstein1: Why don't you try the 3rd question?

OpenStudy (anonymous):